### v1

# Classical and quantum thermodynamics in a non–equilibrium regime: Application to Stirling engine

Abstract
We develop a thermodynamic theory in the non–equilibrium regime by extending a theory based on the dimensionless (DL) minimum work principle previously developed for a thermodynamic system–bath model [S. Koyanagi and Y. Tanimura, J. Chem. Phys. in press (http://arxiv.org/abs/2405.16787)]. Our results are described by non–equilibrium thermodynamic potentials expressed in time-derivative form in terms of extensive and intensive variables. This is made possible through the incorporation of waste heat, which is equivalent to the loss work consumed by the bath, into the definitions of thermodynamic potentials in a non–equilibrium regime. These potentials can be evaluated from the DL non–equilibrium-to-equilibrium minimum work principle, which is derived from the principle of DL minimum work and is equivalent to the second law of thermodynamics. We thus obtain the non–equilibrium Massieu–Planck potentials as entropic potentials and the non–equilibrium Helmholtz–Gibbs potentials as free energies. Our results are verified numerically by simulating a Stirling engine consisting of two isothermal and two thermostatic processes using the quantum Fokker–Planck equations and the classical Kramers equation derived from the thermodynamic system–bath model. We then show that any thermodynamic process can be analyzed using a non–equilibrium work diagram analogous to the equilibrium one for given time-dependent intensive variables. The results can be used to develop efficient heat machines in non–equilibrium regimes.
I Introduction
Ever since Carnot explored the efficiency of heat engines 200 years ago,Carnot (1824) there have been longstanding attempts to study thermodynamics in non–equilibrium regimes, driven by both academic curiosity and practical interest. In particular, recent advances in nanotechnology have led to increased interest in the study of the thermodynamics of microscopic quantum systems. Quan et al. (2006); Esposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Kosloff and Levy (2014); Esposito, Ochoa, and Galperin (2015); Schmidt et al. (2015); Uzdin, Levy, and Kosloff (2015); Whitney (2018); Quan et al. (2007); Maruyama, Nori, and Vedral (2009); Ono et al. (2020); Maruyama, Nori, and Vedral (2009); Talkner and Hänggi (2020a); Cangemi et al. (2020); Strasberg and Winter (2021); Cockrell and Ford (2022); Talkner and Hänggi (2020b); Dann and Kosloff (2023); Ferreri et al. (2023); Sakamoto and Tanimura (2021); Kato and Tanimura (2015, 2016)
As a kinetic theory, such phenomena are typically explained using open quantum dynamics theories based on the system–bath (SB) model.Leggett et al. (1987); Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012) Although thermodynamics is a system-independent theory, it is derived from the presence of a heat bath and is consistent with theories based on the SB model.Talkner and Hänggi (2020b) The main (subsystem) system of the SB model can describe a variety of systems, from a simple two-level system to complex molecular systems with many degrees of freedom, and it is also possible to take the classical limit in the case where the system can be described in phase space.Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012); Tanimura and Kubo (1989); Tanimura (1990, 2014, 2006, 2020) Under the condition of a heat bath with infinite degrees of freedom, the total system irreversibly relaxes toward its equilibrium state in time, and so the second law of thermodynamics is naturally obeyed without the need for the assumption of ergodicity as a dynamical system.Tanimura (2006, 2020, 2014, 2015)
However, to investigate whether the thermodynamic laws are satisfied even in quantum cases where the subsystem and bath are quantum mechanically entangled, it is necessary to treat the bath in a non-Markovian and nonperturbative manner so that the thermal equilibrium state of the total systems satisfies the energy conservation law, including the SB interaction. Therefore, equations of motion derived using the Markovian or rotational wave (or secular) approximation, such as the Lindblad equation and the quantum master equation, can only be applied to high-temperature regions where the subsystem exhibits semiclassical dynamics.Tanimura (2006, 2020, 2014, 2015) Over the past 30 years, several methodologies have been developed to accurately describe the effects of quantum entanglement for condensed systems while maintaining strict energy conservation to satisfy the first law of thermodynamics. Such methods include the hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (1990); Ishizaki and Tanimura (2005); Tanimura (2006, 2020, 2014, 2015) the quasi-adiabatic path integral (QUAPI),Makri (1995); Makri and Makarov (1995a, b); Thorwart, Reimann, and Hänggi (2000); Jadhao and Makri (2008); Makri (2014); Segal, Millis, and Reichman (2010) and multiconfigurational time-dependent Hartree (MCTDH).Meyer, Manthe, and Cederbaum (1990); Manthe, Meyer, and Cederbaum (1992); Wang and Thoss (2007) Among these, the HEOM approach is ideal for thermodynamic investigations, not only because it can perform numerically “exact” dynamic simulations, but also because it can evaluate the change in energy of the heat bath and the SB interaction separately, even for processes far from equilibrium.Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2020, 2021); Song and Shi (2017); Gong et al. (2022); Latune, Pleasance, and Petruccione (2023); Boettcher et al. (2024); Koyanagi and Tanimura (2022a, b, 2024a)
While the difficulty of numerical simulation has been solved, there is a fundamental difference between open quantum dynamics theory, which is based on a first-principles description of kinetic systems from a microscopic perspective, and thermodynamic theory, which is based on a phenomenological description of thermal systems from a macroscopic perspective. For example, in quantum mechanics, observables are defined as expectation values, whereas in thermodynamics, they are described by macroscopic intensive and extensive variables. Furthermore, quantum mechanics is a formalism for time evolution, whereas thermodynamics deals primarily with static and quasi-static states near thermal equilibrium. In this respect, most studies of quantum thermodynamics are merely kinetic simulations of open quantum dynamics, and their relation to thermodynamics has not been studied in depth. Thus, the fundamental difference between microscopic quantum mechanics and macroscopic thermodynamics raises many open questions, such as whether the Carnot limit can be violated in a quantum case or the existence of Maxwell’s demon.
The virtue of thermodynamics lies in its ability to describe macroscopic thermal phenomena resulting from complex microscopic interactions in a system-independent manner as changes in thermodynamic potentials described as interrelated intensive and extensive variables through Legendre transformations. This virtue should be preserved when we develop a quantum thermodynamic theory rather than an open quantum dynamical theory, although in either case, the theory must be specific to the SB model.
It has been found that the Gibbs energy can be obtained directly from kinetic simulations or experiments based on the minimum work principle (or Kelvin–Planck statement), expressed as $W(t)\geq\Delta G(t)$, where $W(t)$ is the work done from the outside to the subsystem by external fields and $\Delta G(t)$ is the change in free energy, by evaluating the work in a quasi-static process,Sakamoto and Tanimura (2020, 2021) which allows us to draw a work diagram corresponding to the $P$$V$ diagram.Koyanagi and Tanimura (2022a, b) However, with this approach, the contributions of temperature $T$ and entropy $S$ in the thermodynamic potential expressed as $TdS$ and $SdT$ cannot be evaluated, because the minimum work principle is defined as an isothermal process $dT=0$.
To overcome this limitation, we developed a thermostatic SB model that is defined by a system coupled to multiple heat baths with different temperatures.Koyanagi and Tanimura (2024a) We then extended the minimum work principle to thermostatic processes in a dimensionless (DL) form (the DL minimum work principle) as ${\tilde{W}}(t)\geq\Delta\Xi(t)$, where ${\tilde{W}}(t)\equiv\beta(t)W(t)$ is the DL (entropic) work, with $\beta(t)\equiv 1/k_{\rm B}T(t)$ (where $k_{\rm B}$ is the Boltzmann constant) being the time-dependent inverse temperature, and $\Delta\Xi$ is the change in the DL Planck potential.Koyanagi and Tanimura (2024a) Not only intensive variables but also extensive variables, which are related by the time-dependent Legendre transformations, were introduced as quantum expectation values of the SB system. The validity of these results has been numerically verified using the low-temperature quantum Fokker–Planck equations in the quantum case and the Kramers equation in the classical case,Ikeda and Tanimura (2019) both developed for the thermodynamic SB model.Koyanagi and Tanimura (2024a) Although these results were restricted to quasi-static cases, the definitions of the thermodynamic potentials, intensive and extensive variables, and Legendre transformations were defined in such a way that they hold for any non–equilibrium process. Taking advantage of this, we extend our thermodynamic theory here to the non–equilibrium regime.
The rest of this paper is organized as follows. In Sec. II, our previous results are summarized with regard to thermodynamics as applied to work in a system–independent manner. We then derive the principle of non–equilibrium DL minimum work to obtain the DL Massieu–Planck potentials and Helmholtz–Gibbs potentials in the non–equilibrium regime. Results are verified in Sec. III by numerical simulations using the SB model. Finally, Sec. IV presents concluding remarks.
II Reflections on work of heat
In our previous paper,Koyanagi and Tanimura (2024a) we presented system–specific thermodynamic laws based on open quantum dynamics theory. Here, we develop the same laws by introducing several thermodynamic “statements” without going into the details of the system, as in traditional thermodynamic theory. By doing so, we clarify the distinctive features of thermodynamic theory that allow it to treat systems in non–equilibrium regimes.
II.1 Laws of thermodynamics as applied to work
We consider a thermodynamic system consisting of a subsystem A and a heat bath B at an inverse temperature $\beta(t)$. The presence of a heat bath temperature can be regarded as a consequence of the zeroth law of thermodynamics.
An important statement of thermodynamics is that thermodynamic systems are described in terms of intensive and extensive variables. In particular, extensive variables that are proportional to the size are essential. This statement, controversial, is sometimes called the fourth law.Oono (2017) However, this is the premise of thermodynamics and we should call the minus first law of thermodynamics.
The energy of the subsystem corresponds to internal energy and is expressed as $U_{\rm A}(t)$, which changes with time owing to changes in the bath temperature. Internal energy is an extensive variable, whereas inverse temperature is an intensive variable. The external perturbation we consider here is expressed as $-x(t)X_{\rm A}(t)$, where $x(t)$ and $X_{\rm A}(t)$ are intensive and extensive variables, respectively. In traditional thermodynamics, $x(t)$ is derived from the Euler relation. We then introduce the total energy, which is related to enthalpy by ${H}_{\rm A}(t)=U_{\rm A}(t)-x(t)X_{\rm A}(t)$, which can also be regarded as the Legendre transformation between internal energy and enthalpy.
Because $U_{\rm A}(t)$ is the conjugate variable of $\beta(t)$, we introduce the DL (or entropic) internal energy denoted by $\tilde{U}_{\rm A}(t)$, where $\tilde{B}(t)\equiv\beta(t)B(t)$ for any variable $B(t)$. Because $\beta(t)$ diverges for $T\rightarrow 0$, states with $T=0$ do not exist, which corresponds to the third law of thermodynamics.
From the time derivative of the DL internal energy $\tilde{U}_{\rm A}(t)$, we have
$\frac{d\tilde{U}_{\rm A}(t)}{dt}=\frac{d\tilde{W}^{ext}_{\rm A}(t)}{dt}+\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt},$ (1)
where
$\frac{d\tilde{W}^{ext}_{\rm A}(t)}{dt}=U_{\rm A}(t)\frac{d\beta(t)}{dt}+\tilde{x}(t)\frac{dX_{\rm A}(t)}{dt}$ (2)
and
$\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}\equiv\beta(t)\frac{dU_{\rm A}(t)}{dt}-\tilde{x}(t)\frac{dX_{\rm A}(t)}{dt}$ (3)
are the DL (or entropic) extensive work defined as being done from the outside and DL (or entropic) extensive heat, respectively. For the enthalpy, we have
$\frac{d\tilde{H}_{\rm A}(t)}{dt}=\frac{d\tilde{W}_{\rm A}^{int}(t)}{dt}+\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt},$ (4)
where
$\frac{d\tilde{W}_{\rm A}^{int}(t)}{dt}\equiv U_{\rm A}(t)\frac{d\beta(t)}{dt}-X_{\rm A}(t)\frac{d\tilde{x}(t)}{dt}$ (5)
is the intensive work. Equations (1) and (4) correspond to the first law of thermodynamics.
From the definitions, these intensive and extensive variables satisfy the time-dependent Legendre (TDL) transformation expressed as
$\frac{d\tilde{W}_{\rm A}^{int}(t)}{dt}=\frac{d\tilde{W}^{ext}_{\rm A}(t)}{dt}-\frac{d}{dt}\left[\tilde{x}(t)X_{\rm A}(t)\right].$ (6)
Equation (1) is also expressed in the form of the TDL transformation as
$\frac{d\tilde{W}_{\rm A}^{ext}(t)}{dt}=-\frac{d\tilde{Q}^{ext}_{\rm A}(t)}{dt}+\frac{d}{dt}\left[\beta(t)U_{\rm A}(t)\right].$ (7)
The intensive heat $\tilde{Q}_{\rm A}^{int}(t)$ can also be defined from $\tilde{Q}_{\rm A}^{ext}(t)$ using the TDL [Eq. (46) in Appendix A]. While the above equations, including the first law of thermodynamics, hold for any non–equilibrium processes described by intensive variables $\beta(t)$ and $\tilde{x}(t)$ and extensive variables $U_{\rm A}(t)$ and $X_{\rm A}(t)$ which are the state variables, as they are kinematic observables at time $t$, work and heat are not state variables.
As the second law of thermodynamics, we adopt the DL minimum work principle for the subsystem from one equilibrium state to another, which extends the Kelvin-Planck statement for isothermal processes to thermostatic processes. For work defined as being done from the outside to the subsystem, we have
$\tilde{W}_{\rm A}^{ext}\geq-\Delta\Phi_{\rm A}^{\rm qst},$ (8)
where $\Phi_{\rm A}^{\rm qst}$ is the DL Massieu potential Massieu (1869) and equality holds under quasi-static changes of natural variables $\beta(t)$ and $X_{\rm A}(t)$ as $\Phi_{\rm A}^{\rm qst}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]=\tilde{W}_{\rm A}^{ext}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]$. The above inequality states that for any process occurring between two states, the work done from the outside is minimized if the process is quasistatic (or reversible), because no energy is dissipated in the heat bath. Note that if we define work as being done from the subsystem to outside, the inequality in Eq.(8) is reversed and the relationship is called the maximum work principle (to the outside). Using the TDL transformation (6), we also have the Planck potential Planck (1922) in terms of natural variables expressed as $\Xi_{\rm A}^{\rm qst}[\beta^{\rm qst},\tilde{x}^{\rm qst}]$ (see Appendix A).
For the DL extensive heat $\tilde{Q}_{\rm A}^{ext}(t)$, we have
$\tilde{Q}_{\rm A}^{ext}\leq\Delta\Lambda_{\rm A}^{\rm qst},$ (9)
which corresponds to the principle of maximum entropy, where $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$ is the entropic potential. Because both natural variables are extensive, this potential is fundamental and we call it Massieu entropy (M-entropy). Using the TDL tranformation Eq. (46), we also have Planck entropy (P-entropy) defined by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$ from $\tilde{Q}_{\rm A}^{int}(t)$ (see Appendix A). Similar to the entropy obtained from the partition function in statistical physics, this entropy is a function of the extensive variable $\tilde{x}^{\rm qst}$.
Here, as the conjugate variable for $\beta(t)$, we chose the internal energy $U_{\rm A}^{\rm qst}$, which is considered a fundamental variable in thermodynamics, but enthalpy $H_{\rm A}^{\rm qst}$ could also be chosen.Koyanagi and Tanimura (2024a) In such case, there are also two entropies, one involving both extensive variables (Clasius entropy or C-entropy) and one for intensive variable with respect to external forces (Boltzmann entropy or B-entropy). The B-entropy and M-entopy are related to the others by the Legendre transformations between $U_{\rm A}^{\rm qst}$ and $H_{\rm A}^{\rm qst}$ and the values do not change, although the former inculdes one intensive variable whereas the later includes extensive variables only. However, there is no explicit relationship between C-entropy and M-entropy, because the Legendre transformation between two entropy in the $U_{\rm A}^{\rm qst}$ representation and the $H_{\rm A}^{\rm qst}$ representation are different. (See Appendix D in Ref. Koyanagi and Tanimura, 2024a). In any case, the Massieu–Planck potentials are entropic potentials.Guggenheim (1986)
The total differential form of the Helmholtz–Gibbs potentials can be derived from the Massieu–Planck potentials using the definitions $F^{\rm qst}_{\rm A}(t)=-\Phi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$ and $G^{\rm qst}_{\rm A}(t)=-\Xi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$. The results are summarized and presented in Table 6 in Appendix A.
II.2 Thermodynamic potentials in a non–equilibrium regime
II.2.1 The DL non–equilibrium minimum work principle
On the basis of Eq. (8), we can define thermodynamic potentials applicable to the non–equilibrium regime. Consider a non–equilibrium state $A$ and equilibrium state $n$. The DL non–equilibrium-to-equilibrium minimum work principle for $A\rightarrow n$ is expressed as (see Appendix B)
$(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}\geq-(\Delta\Phi_{A}^{\rm neq})_{A\rightarrow n},$ (10)
where $\Phi_{\rm A}^{\rm neq}(t)$ is the non–equilibrium Massieu potential. This inequality indicates that the path from non–equilibrium to equilibrium has a lower bound on work. In engineering, the effective energy from non-equilibrium to equilibrium is referred to as exergy.Rant (1956); Rosen and Dincer (2001) The non–equilibrium thermodynamic potential we introduced here can be regarded as a generalization of it.
The difference between $(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}$ and $-(\Delta\Phi_{A}^{\rm neq})_{{A}\rightarrow n}$ corresponds to the DL heat generated as loss work, denoted by $-(Q_{\rm A}^{\rm wst})_{{A}\rightarrow n}$, which we call DL waste heat:
$-(\tilde{Q}_{\rm A}^{\rm wst})_{{A}\rightarrow n}=(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}+(\Delta\Phi_{\rm A}^{\rm neq})_{{A}\rightarrow n}\geq 0.$ (11)
Because the DL waste heat between two non–equilibrium states $A$ and $B$ is evaluated as $(\tilde{Q}_{\rm A}^{\rm wst})_{{A}\rightarrow{B}}=(\tilde{Q}_{\rm A}^{\rm wst})_{{A}\rightarrow n}-(\tilde{Q}_{\rm A}^{\rm wst})_{B\rightarrow n}$, we the waste heat for any non–equilibrium process at times $t$ and $t+dt$ (where $dt$ is an infinitesimal time) in time-derivative form:
$-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}=\frac{d\tilde{W}_{\rm A}^{ext}(t)}{dt}+\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}.$ (12)
As shown in Appendix B, the waste heat is nonpositive, i.e., $d\tilde{Q}_{\rm A}^{\rm wst}(t)/dt\leq 0$. Thus, we have the inequality $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow B}^{\rm min}\geq-(\Delta\Phi_{\rm A}^{\rm neq})_{\rm A\rightarrow B}$, or
$\frac{d\tilde{W}_{\rm A}^{ext}(t)}{dt}\geq-\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt},$ (13)
which we call the DL non–equilibrium minimum work principle. While $\tilde{W}_{\rm A}^{ext}(t)$ and $\tilde{Q}_{\rm A}^{\rm wst}(t)$ are not state variables because they depend on a path, $\Phi_{\rm A}^{\rm neq}(t)$ is a state variable defined by the non–equilibrium-to-equilibrium minimum work path. When state $B$ is on the non–equilibrium-to-equilibrium minimum work path from $A$ to $n$, equality holds in (13).
Using Eqs. (2) and (12), we obtain the time-derivative form of the non–equilibrium Massieu potential as
$\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}=-U_{\rm A}^{\rm neq}\frac{d\beta(t)}{dt}-\tilde{x}(t)\frac{d{X}_{\rm A}^{\rm neq}(t)}{dt}-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}.$ (14)
The convexity of DL potentials is discussed on the basis of the SB model in Appendix C. Thus, we can regard these as thermodynamic potentials.
Using Eqs. (6), we have the non–equilibrium Planck potential expressed as
$\frac{d\Xi_{\rm A}^{\rm neq}(t)}{dt}=-U_{\rm A}^{\rm neq}\frac{d\beta(t)}{dt}+{X}_{\rm A}^{\rm neq}(t)\frac{d\tilde{x}(t)}{dt}-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}.$ (15)
The M-entropy $\Lambda_{\rm A}^{\rm neq}(t)$ and P-entropy $\Gamma_{\rm A}^{\rm neq}(t)$ can be evaluated from the TDL transformations (6) and  (7). The non–equilibrium potentials satisfy the Legendre transformations:
$\Xi_{\rm A}^{\rm neq}(t)=\Phi_{\rm A}^{\rm neq}(t)+\tilde{x}(t){X}_{\rm A}^{\rm neq}(t)$ (16)
and
$\Lambda_{\rm A}^{\rm neq}(t)=\Phi_{\rm A}^{\rm neq}(t)+\beta(t)U_{\rm A}^{\rm neq}(t).$ (17)
The time-derivative forms of the DL non–equilibrium entropic potentials are summarized in Table 1. Note that while the entropic potentials, the intensive variables and non–equilbrium extensive variables are state variables, the time–deribative of them including waste heat are not state variables.
As shown in Appendix D, the non–equilibrium Massieu potential is always smaller than the quasi-static Massieu potential. This indicates that the non–equilibrium Massieu potential is minimum when the state is equilibrium. In the quasi-static case, Table  1 is reduced to Table 5 (Appendix A). Thus, we can regard these potentials as extensions of the thermodynamic potentials to a non–equilibrium regime.
From Eqs. (12) and (17), the waste heat current is expressed using the non–equilibrium M-entropy as
$\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}=\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}-\frac{d\Lambda_{\rm A}^{\rm neq}(t)}{dt}.$ (18)
Table 1: Time-derivative forms of the non–equilibrium (neq) Massieu–Planck potentials as functions of the intensive variables $\beta(t)$ and $\tilde{x}(t)$ and the extensive variables $U_{\rm A}^{\rm neq}(t)$ and $X_{\rm A}^{\rm neq}(t)$. Of the DL entropies, the commonly used one, which we call the Massieu entropy (M-entropy), involves only extensive variables and is denoted by $\Lambda_{\rm A}^{\rm neq}[U_{\rm A}^{\rm neq}(t),{X}_{\rm A}^{\rm neq}(t)]$, whereas the less widely used one, which we call the Planck entropy (P-entropy), is denoted by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}]$. Each potential is related to the others by the Legendre transformations shown in the final column.
neq DL Thermodynamic pot. Differential Form Natural var. Legendre Transformation
Massieu $\frac{d}{dt}\Phi_{\rm A}^{\rm neq}=-U_{\rm A}^{\rm neq}\frac{d}{dt}\beta-\tilde{x}\frac{d}{dt}{X}_{\rm A}^{\rm neq}-\frac{d}{dt}\tilde{Q}^{\rm wst}_{\rm A}$ $\beta,{X}_{\rm A}^{\rm neq}$cdots
Planck $\frac{d}{dt}\Xi^{\rm neq}_{\rm A}=-U_{\rm A}^{\rm neq}\frac{d}{dt}\beta+{X}_{\rm A}^{\rm neq}\frac{d}{dt}\tilde{x}-\frac{d}{dt}\tilde{Q}^{\rm wst}_{\rm A}$ $\beta,\tilde{x}$ $\Xi_{\rm A}^{\rm neq}=\Phi_{\rm A}^{\rm neq}+\tilde{x}{X}_{\rm A}^{\rm neq}$
M-Entropy $\frac{d}{dt}\Lambda_{\rm A}^{\rm neq}=\beta\frac{d}{dt}U_{\rm A}^{\rm neq}-\tilde{x}\frac{d}{dt}X_{\rm A}^{\rm neq}-\frac{d}{dt}\tilde{Q}_{\rm A}^{\rm wst}$ $U_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $\Lambda_{\rm A}^{\rm neq}=\Phi_{\rm A}^{\rm neq}+\beta U_{\rm A}^{\rm neq}$
P-Entropy $\frac{d}{dt}\Gamma_{\rm A}^{\rm neq}=\beta\frac{d}{dt}U_{\rm A}^{\rm neq}+X_{\rm A}^{\rm neq}\frac{d}{dt}\tilde{x}-\frac{d}{dt}\tilde{Q}_{\rm A}^{\rm wst}$ $U_{\rm A}^{\rm neq},\tilde{x}$ $\Gamma_{\rm A}^{\rm neq}=\Lambda_{\rm A}^{\rm neq}+\tilde{x}{X}_{\rm A}^{\rm neq}$
II.2.2 Nonequilibrium Helmholtz–Gibbs potentials
Table 2: Time-derivative forms of the non–equilibrium (neq) Helmholtz–Gibbs potentials as functions of the intensive variables $T(t)$ and $x(t)$ and the extensive variables $S_{\rm A}^{\rm neq}(t)$ and $X_{\rm A}^{\rm neq}(t)$, which are interrelated through the Legendre transformations shown in the final column. Because heat is always lost in non–equilibrium processes, the waste heat $Q^{\rm wst}(t)$ appears in the equations.
neq Thermodynamic pot. Differential Form Natural var. Legendre Transformation
Helmholtz Energy $\frac{d}{dt}F^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q^{\rm wst}$ $T,x$
Gibbs Energy $\frac{d}{dt}G^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q^{\rm wst}$ $T,X_{\rm A}^{\rm neq}$ $G^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
Internal Energy $\frac{d}{dt}U^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q^{\rm wst}$ $S_{\rm A}^{\rm neq},x$ $U^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}+TS^{\rm neq}_{\rm A}$
Enthalpy $\frac{d}{dt}H^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q^{\rm wst}$ $S_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $H^{\rm neq}_{\rm A}=U^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
We introduce non–equilibrium Helmholtz and Gibbs energies defined as $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta(t)$ and $G^{\rm neq}_{\rm A}(t)=-\Xi_{\rm A}^{\rm neq}(t)/\beta(t)$. Because $d\beta(t)/dt=-(1/k_{\rm B}T^{2}(t))dT(t)/dt$, we obtain the time-derivative forms of these from Eqs. (14) and (15) as
$\frac{dF_{\rm A}^{\rm neq}(t)}{dt}=-S_{\rm A}^{\rm neq}(t)\frac{dT(t)}{dt}+x(t)\frac{dX_{\rm A}^{\rm neq}(t)}{dt}+\frac{dQ_{\rm A}^{\rm wst}(t)}{dt}$ (19)
and
$\frac{dG_{\rm A}^{\rm neq}(t)}{dt}=-S_{\rm A}^{\rm neq}(t)\frac{dT(t)}{dt}+X_{\rm A}^{\rm neq}(t)\frac{dx(t)}{dt}+\frac{dQ_{\rm A}^{\rm wst}(t)}{dt},$ (20)
where $S_{\rm A}^{\rm neq}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm neq}(t)$ is a non–equilibrium entropy and we have introduced the waste heat current as
$\frac{dQ_{\rm A}^{\rm wst}(t)}{dt}=\frac{1}{\beta(t)}\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}.$ (21)
From Eq. (16), we obtain the TDL transformation between the non–equilibrium Gibbs and Helmholtz potentials as
$F_{\rm A}^{\rm neq}(t)=G_{\rm A}^{\rm neq}(t)+x(t)X_{\rm A}^{\rm neq}(t).$ (22)
Solving Eq. (17) for $U_{\rm A}^{\rm neq}(t)$, we obtain the following TDL transformations:
$U_{\rm A}^{\rm neq}(t)=F_{\rm A}^{\rm neq}(t)+T(t)S_{\rm A}^{\rm neq}(t).$ (23)
From the above, we obtain the time-derivative expressions for the enthalpy and internal energy. The results are summarized in Table 2. These results reduce to those in Table 6 (Appendix A) in the quasi-static case. Thus, we can regard these potentials as extensions of the thermodynamic potentials to a non–equilibrium regime.
Dividing both sides of Eq. (3) by $\beta(t)$ yields the first law of thermodynamics expressed for the internal energy as
$\frac{dU_{\rm A}^{\rm neq}(t)}{dt}=\frac{dW_{\rm A}^{ext}(t)}{dt}+\frac{dQ^{ext}_{\rm A}(t)}{dt},$ (24)
where
$\frac{dQ_{\rm A}^{ext}(t)}{dt}=\frac{1}{\beta(t)}\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}$ (25)
is the extensive heat current. From Eqs. (18) and (25), we can evaluate the DL waste heat current as
$-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}=\frac{1}{k_{\rm B}}\left\{\frac{dS_{\rm A}^{\rm neq}(t)}{dt}-\frac{1}{T(t)}\frac{dQ_{\rm A}^{ext}(t)}{dt}\right\}.$ (26)
In the isothermal case, by integrating both sides between the two equilibrium states over time $t$, we obtain
$-\tilde{Q}_{\rm A}^{\rm wst}=\frac{1}{k_{\rm B}}\left\{\Delta S_{\rm A}^{\rm qst}-\frac{Q_{\rm A}^{ext}}{T}\right\},$ (27)
which indicates that the DL waste heat is equivalent to the entropy production evaluated for a heat engine.Sakamoto and Tanimura (2020); Koyanagi and Tanimura (2022a)
The non–equilibrium Gibbs energy satisfies
$S^{\rm neq}_{\rm A}(t)=-\left(\frac{\partial G^{\rm neq}_{\rm A}}{\partial T(t)}\right)_{x(t),\tilde{Q}^{\rm wst}},$ (28)
$X_{\rm A}(t)=-\left(\frac{\partial G_{\rm A}^{\rm neq}}{\partial x(t)}\right)_{T(t),\tilde{Q}^{\rm wst}},$ (29)
and
$H_{\rm A}(t)=-T^{2}(t)\frac{\partial}{\partial T(t)}\left(\frac{G^{\rm neq}_{\rm A}(t)}{T(t)}\right)_{x(t),\tilde{Q}^{\rm wst}}$ (30)
Equation (30) is an extension of the Gibbs–Helmholtz relation to a non–equilibrium regime.
III Numerical Demonstration
Although the results presented in Tables 1 and  2 hold for any non–equilibrium system consisting of subsystem and bath, the extensive variables and waste heat that appear in these relationships are system-specific, and there is no general theory on the basis of which non–equilibrium thermodynamic potentials can be obtained. However, it is possible to evaluate them as functions of the intensive and extensive variables using an optimization algorithm. Here, as a demonstration, we evaluate non–equilibrium thermodynamic potentials numerically using the thermodynamic SB model.Koyanagi and Tanimura (2024a)
III.1 Thermodynamic system–bath model
We employ the Ullersma–Caldeira–Leggett (or Brownian) model,Ullersma (1966a, b); Caldeira and Leggett (1983); Grabert, Schramm, and Ingold (1988); Weiss (2012); Tanimura and Wolynes (1991, 1992) which is ideal for thermodynamic simulations because the subsystem and bath are well defined, and rigorous numerical solutions can be obtained in both classical and quantum cases under any time-dependent external perturbation. Many of the favorable features for thermodynamic investigations arise from the presence of a counter term, which allows us to include the contribution of the SB interaction in the bath.Tanimura (2015, 2006, 2020); Tanimura and Wolynes (1991, 1992); Kato and Tanimura (2013) By introducing multiple heat baths at different temperatures controlled by time-dependent SB coupling functions, we can investigate isothermal, isentropic, thermostatic, and entropic processes. The total Hamiltonian for isothermal and thermostatic processes is written asKoyanagi and Tanimura (2024a)
$\hat{H}_{\rm tot}(t)=\hat{H}_{A}^{0}+\hat{H}_{\rm A}^{\prime}(t)+\sum_{k=0}^{N}\hat{H}_{\rm IB}^{k}(t),$ (31)
where
$\hat{H}_{\rm A}^{0}=\frac{\hat{p}^{2}}{2m}+U(\hat{q})$ (32)
is the unperturbed Hamiltonian of the subsystem with mass $m$ and potential $U(\hat{q})$ described by momentum $\hat{p}$ and position $\hat{q}$. The internal energy is then evaluated as $U_{\rm A}^{\rm neq}(t)=\mathrm{tr}\{\hat{H}_{\rm A}^{0}\hat{\rho}_{\rm A}(t)\}$, where $\hat{\rho}_{\rm A}(t)$ is the reduced density operator of the subsystem. The external perturbation is expressed as $\hat{H}_{\rm A}^{\prime}(t)\equiv-x(t)\hat{X}_{\rm A}$, where $\hat{X}_{\rm A}$ is an operator of the subsystem coordinate [i.e., $\hat{X}_{\rm A}(\hat{q})$] and $x(t)$ is regarded as the thermodynamic intensive variable. The extensive variable is evaluated as $X_{\rm A}^{\rm neq}(t)=\mathrm{tr}\{\hat{X}_{\rm A}\hat{\rho}_{\rm A}(t)\}$.
Although the conventional SB model has been limited to the investigation of isothermal processes at constant temperature, we can extend it to describe thermostatic processes in which temperature varies with time by introducing $N$ independent heat baths, each in the thermal equilibrium state at the inverse temperature $\beta_{k}\equiv 1/k_{\rm B}T_{k}$ connected to or disconnected from subsystem A using the window function $\xi_{k}(t)$.Koyanagi and Tanimura (2024a) The $k$th bath Hamiltonian is expressed as an ensemble of harmonic oscillators and is given by
$\hat{H}_{\rm IB}^{k}(t)\equiv\sum_{j}\left\{\frac{(\hat{p}^{k}_{j})^{2}}{2m_{j}}+\frac{m_{j}^{k}(\omega_{j}^{k})^{2}}{2}\left[\hat{x}_{j}^{k}-\frac{c_{j}^{k}A_{k}\xi_{k}(t)\hat{q}}{m_{j}^{k}(\omega_{j}^{k})^{2}}\right]^{2}\right\},$ (33)
where the momentum, position, mass, and frequency of the $j$th bath oscillator are given by $\hat{p}_{j}^{k}$, $\hat{x}_{j}^{k}$, $m_{j}^{k}$, and $\omega_{j}^{k}$, respectively. Here, we consider the situation where $N$ independent heat baths, each in the thermal equilibrium state $\exp(-\beta_{k}\hat{H}_{\rm I+B}^{k})$ at the inverse temperature $\beta_{k}\equiv 1/k_{\rm B}T_{k}$, are connected or disconnected to subsystem A according to a control function $\xi_{k}(t)$. The bath temperature is effectively expressed as
$T(t)=\sum_{k=1}^{N}T_{k}\xi_{k}(t),$ (34)
or the inverse temperature as $\beta(t)=[k_{\rm B}T(t)]^{-1}$. As a spectral distribution function for the $k$th bath $J^{k}(\omega)\equiv\sum_{j}{\hbar(c_{j}^{k}A_{k})^{2}}/({2m_{j}^{k})\omega_{j}^{k}}\delta(\omega-\omega_{j}^{k})$, we consider the Ohmic case described by
$J^{k}(\omega)=\frac{\hbar A_{k}^{2}\omega}{\pi}$ (35)
and assume that the time scale of the quantum thermal fluctuations $\beta(t)\hbar/2\pi$ is shorter than the time scale of the external perturbations $\beta(t)$ and $x(t)$. This allows us to extend the low-temperature quantum Fokker–Planck equations (LT-QFPE)Ikeda and Tanimura (2019) for the set of Wigner distribution functions $W_{\vec{n}}(p,q;t)$, where ${\vec{n}}$ is the index of hierarchy member, and the Kramers equation for phase space distribution function $W(p,q;t)$ to apply the thermostatic case by introducing time-dependent Matsubara frequencies $\nu(t)=1/\hbar\beta(t)$. Expressions for these equations are presented in Ref. Koyanagi and Tanimura, 2024a. The source code for thermodynamic HEOM in phase space representation and thermodynamic Kramers equation are provided as a separated paper.Koyanagi and Tanimura (2024b).
It should be noted that in the non-Ohmic case, the temperature may change even on the time scale on which the noise correlations are defined. This makes it difficult to apply the fluctuation–dissipation theorem to characterize the noise, and thus a simple replacement $\beta\rightarrow\beta(t)$ is not allowed.
III.2 DL work and extensive variables
We express the solution for the reduced density elements under any $x(t)$ in the Wigner representation using the zeroth member of the hierarchical Wigner functions as $W(p,q,t)\equiv W_{\vec{0}}(p,q;t)$. In the classical limit $\hbar\rightarrow 0$, ${W}(q,p;t)$ corresponds to the classical distribution function. We use this to define the change in DL intensive work over time corresponding to power and heat flow as follows:
$\frac{d\tilde{W}_{\rm A}^{int}(t)}{dt}=U_{\rm A}^{\rm neq}(t)\frac{d\beta(t)}{dt}-X_{\rm A}^{\rm neq}(t)\frac{d\tilde{x}(t)}{dt},$ (36)
where the extensive variables in the Wigner representation at time $t$ are expressed as
$X_{\rm A}^{\rm neq}(t)=\mathrm{tr}_{\rm A}\{X_{\rm A}(q){W}(p,q;t)\}$ (37)
and
$U_{\rm A}^{\rm neq}(t)=\mathrm{tr}_{\rm A}\left\{\left[\frac{p^{2}}{2m}+U(q)\right]W(p,q;t)\right\}.$ (38)
From the definition (2) of extensive work and the TDL transformations (6) and (7), the extensive heat current is evaluated as
$\frac{dQ_{\rm A}^{ext}(t)}{dt}=\\ \mathrm{tr}_{\rm A}\left\{\left[\frac{p^{2}}{2m}+U(q)-x(t)X_{\rm A}(q)\right]\frac{\partial W(p,q;t)}{\partial t}\right\}.$ (39)
III.3 Evaluation of the non–equilibrium Planck potential
Figure 1: Schematic of the evaluation of the non–equilibrium Planck potential $\Xi_{\rm A}^{\rm neq}(t)$ at $t_{k}$ and $t_{k+1}$. A non–equilibrium process driven by the intensive variables $\beta(t)$ and $\tilde{x}(t)$ is described as the non–equilibrium distribution $W(p,q;t)$ expressed by the black arrows. The equilibrium states for different sets of $\beta^{\rm qst}_{n_{k}}$ and $\tilde{x}_{n_{k}}^{\rm qst}$ are expressed by the red dashed line. For each state $k$ to the equilibrium state $n_{k}$, we minimize $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}+(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ by choosing the trial functions $\beta_{k}^{\rm trial}(t^{\prime})$ and $\tilde{x}_{k}^{\rm trial}(t^{\prime})$, where $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is the DL minimum intensive work for $k\rightarrow n_{k}$ and $(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ is the quasi-static Planck potential at the state $n_{k}$. We can then evaluate the non–equilibrium Planck potential at the state $k$ as $(\Xi_{\rm A}^{\rm neq})_{k}=(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}+(\Xi_{\rm A}^{\rm neq})_{n_{k}}$.
From Eq. (51), we obtain the expression for the non–equilibrium Planck potential $(\Xi_{\rm A}^{\rm neq})_{k}=\Xi_{\rm A}^{\rm neq}[\beta(t_{k}),\tilde{x}(t_{k})]$ of the non–equilibrium-to-equilibrium process $k\rightarrow n_{k}$ as
$(\Xi_{\rm A}^{\rm neq})_{k}=(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}+(\Xi_{\rm A}^{\rm qst})_{n_{k}},$ (40)
where $n_{k}$ represents an equilibrium state, $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is the DL minimum intensive work for $k\rightarrow n_{k}$, and $(\Xi_{\rm A}^{\rm qst})_{n_{k}}$ is the quasi-static Planck potential at state $n_{k}$. As shown in Eq. (55), because the right-hand side of Eq. (40) is independent of the choice of $n_{k}$, to evaluate $(\Xi_{\rm A}^{\rm neq})_{k}$, we do not specify $n_{k}$ to minimize $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}+(\Xi_{\rm A}^{\rm qst})_{n_{k}}$.
To perform numerical calculations, we express Eq. (40) with Eq. (15) in terms of trial functions $\beta^{\rm trial}(t_{k},t^{\prime})$ and $\tilde{x}^{\rm trial}(t_{k},t^{\prime})$ as
$Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),x^{\rm trial}(t)]=\Xi^{\rm qst}_{\rm A}(\beta^{\rm qst}_{n_{k}},x^{\rm qst}_{n_{k}})$
[…]
where $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),\tilde{x}^{\rm trial}(t)]$ is the target function to be minimized and $\Xi_{\rm A}^{\rm qst}(\beta_{n_{k}}^{\rm qst},\tilde{x}_{n_{k}}^{\rm qst})$ is the quasi-static Planck potential at $n_{k}$, which is also evaluated from the LT-QFPE and Kramers equation. The DL trail function $\tilde{x}_{k}^{\rm trial}(t)$ in Eq.(III.3) is evaluated from $x_{k}^{\rm trial}(t)$ as $\tilde{x}_{k}^{\rm trial}(t)=\beta_{k}^{\rm trial}(t)x_{k}^{\rm trial}(t)$. To reduce the computational cost, we assume that the optimized trial functions become constant after the characteristic time for the equilibration $\Delta t$, as $\beta^{\rm qst}_{n_{k}}=\beta_{k}^{\rm trial}(t_{k}+\Delta t)$ and $x_{n_{k}}^{\rm qst}=x_{k}^{\rm trial}(t_{k}+\Delta t)$.
As the trial functions for time $t^{\prime}>t_{k}$, we chose the ${N_{\beta}}$th and ${N_{x}}$th Taylor expansion forms expressed as
$\beta_{k}^{\rm trial}(t^{\prime})=\sum_{n=0}^{N_{\beta}}\beta_{k}^{(n)}(t^{\prime}-t_{k})^{n},$ (42)
and
$x_{k}^{\rm trial}(t^{\prime})=\sum_{n=0}^{N_{x}}x_{k}^{(n)}(t^{\prime}-t_{k})^{n},$ (43)
where $\beta_{k}^{(n)}$ and $x_{k}^{(n)}$ are the $n$th-order Taylor coefficients. Thus, the functional minimization of $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),\tilde{x}^{\rm trial}(t)]$ becomes a multivariable function minimization for $\beta_{k}^{(n)}$ and $x_{k}^{(n)}$.
Then, $\Xi_{\rm A}^{\rm neq}(t)$ is evaluated as follows (see Fig. 1). First, we perform the non–equilibrium simulations for given $\beta(t)$ and $\tilde{x}(t)$ to obtain $W_{\vec{n}}(p,q;t)$ for each $t_{k}$. The minimum work from the non–equilibrium state $k$ to the equilibrium state $n_{k}$ expressed as $(\tilde{W}_{\rm A}^{int})_{k\rightarrow n_{k}}^{\rm min}$ is then evaluated using an optimization algorithm for $\beta_{k}^{\rm trial}(t^{\prime})$ and $\tilde{x}_{k}^{\rm trial}(t^{\prime})$. From Eq. (40), we set this value as $(\Xi_{\rm A}^{\rm neq})_{k}$. By repeating this procedure for different values of $k$, we obtain $\Xi_{\rm A}^{\rm neq}(t_{k})$ at each step. The other potentials can be evaluated from the TDL transformations.
III.4 Quantum and classical Stirling engine
The non–equilibrium thermodynamic potentials are state variables as functions of the non–equilibrium intensive and extensive variables, which are also state variables, while $Q_{\rm A}^{\rm wst}$ is not, and, in the quasi-static limit, they reduce to conventional thermodynamic potentials. These non–equilibrium potentials are useful because they can be used to analyze thermodynamic processes using work diagrams, as in the case of equilibrium thermodynamics. By finding a thermodynamic process to minimize waste heat, we can construct a heat machine with maximum efficiency under non–equilibrium conditions, although the efficiency is lower than the Carnot limit.Koyanagi and Tanimura (2022a, b)
Here, we demonstrate how to evaluate non–equilibrium potentials for a case in which $\beta(t)$ and $x(t)$ are specified. For this purpose, we consider the classical/quantum Stirling engine (CSE/QSE) consisting of four steps: (i) a hot isothermal process, (ii) a transition from a hot to cold thermostatic process, (iii) a cold isothermal process, and (iv) a transition from a cold to hot thermostatic process for an anharmonic potential system. Since the purpose of the calculation is to illustrate how the potentials are calculated, we fix the changes in $\beta(t)$ and $x(t)$ as shown in Fig. 2 with $\tau=20$ and do not optimize them to improve their efficiency.
III.4.1 Simulation details
Figure 2: Time profiles of (a) the electric field $x(t)$ and (b) the inverse temperature $\beta(t)$ for the Stirling engine. The red, green, blue, and orange lines represent (i) hot isothermal, (ii) isoelectric (from hot to cold), (iii) cold isothermal, and (iv) isoelectric (from cold to hot) processes, respectively. Time $t$ is normalized by the period of the cycle $\tau$.
We performed simulations using the anharmonic potential system employed in our previous study.Koyanagi and Tanimura (2024a) Thus, we considered a quartic anharmonic potential with the external interaction described as $X_{\rm A}(\hat{q})=\hat{q}$. The potential function in Eq. (32) is expressed as
$U(\hat{q})=U_{2}\hat{q}^{2}+U_{3}\hat{q}^{3}+U_{4}\hat{q}^{4},$ (44)
where the constants are given by $U_{2}=0.1$, $U_{3}=0.02$, and $U_{4}=0.05$. Numerical calculations were carried out to integrate the thermodynamic HEOM in the quantum cases and the thermodynamic Kramers equation in the classical cases. The detailed conditions for the numerical calculations, including the working parameters, are presented in Table 3 and in Ref. Koyanagi and Tanimura, 2024a
Table 3: Parameter values used for the simulations of the Stirling engine. Here, $dx$ and $dp$ are the mesh sizes for position and momentum, respectively, in Wigner space. The integers $N$ and $K$ are the cutoff numbers used in the LT-FPE.
$A$ $N$ $K$ $dx$ $dp$
Classical 0.50.50.50.51.51.51.51.5 $\cdots$ $\cdots$ $0.25$ $0.25$
Quantum $0.5$ $6$ $2$ $0.3$ $0.4$
$1.0$ $7$ $2$ $0.3$ $0.5$
$1.5$ $8$ $2$ $0.3$ $0.6$
To set the trial functions defined in Eqs. (42) and (43), we chose $N_{\beta}=N_{x}=5$. Then, using the Nelder–Mead method, we minimized $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),x^{\rm trial}(t)]$ in Eq. (III.3) with a cutoff time $\Delta t=1.0$. We evaluated $\Xi_{\rm A}^{\rm neq}(t)$ every 20 steps. The value of the quasi-static Planck potential $\Xi_{\rm A}^{\rm qst}(\beta^{\rm qst},\tilde{x}^{\rm qst})$ in Eq. (III.3) was obtained by integrating the thermodynamic LT-QFPE and the Kramers equation.
III.4.2 Results
Figure 3: $x(t)$$X_{\rm A}^{\rm neq}(t)$ diagrams for the Stirling engine in the classical case (1, left column) and quantum case (2, right column) for (a) $A=0.5$ (weak), (b) $1.0$ (intermediate), and (c) $1.5$ (strong) SB coupling strengths. In each plot, the four curves (or lines) represent (i) hot isothermal (red), (ii) from hot to cold thermostatic (green), (iii) cold isothermal (blue), and (iv) from cold to hot thermostatic (orange) processes, respectively. The processes evolve in a counterclockwise fashion over time in a heat engine, whereas they evolve in a clockwise fashion over time in a refrigerator.
Figure 4: $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams for the Stirling engine in the classical case (1, left column) and quantum case (2, right column) for (a) $A=0.5$ (weak), (b) $1.0$ (intermediate), and (c) $1.5$ (strong) SB coupling strength, respectively. Each cycle starts with the red arrow, and the four curves represent (i) hot isothermal (red), (ii) from hot to cold thermostatic (green), (iii) cold isothermal (blue), and (iv) from cold to hot thermostatic (orange) processes, respectively.
Table 4: Work performed in one cycle for the classical and quantum cases for different SB coupling strengths.
$A$ $W$ (classical) $W$ (quantum)
$0.5$ (weak) $-1.44\times 10^{-2}$ $-1.07\times 10^{-2}$
$1.0$ (intermediate) $-5.77\times 10^{-3}$ $-3.09\times 10^{-3}$
$1.5$ (strong) $1.01\times 10^{-2}$ $1.10\times 10^{-2}$
The extensive variables $X_{\rm A}^{\rm neq}(t)$ and $U_{\rm A}^{\rm neq}(t)$ were obtained from Eqs. (37) and (38), respectively. The non–equilibrium entropy expressed as $S_{\rm A}^{\rm neq}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm neq}(t)$ was then obtained from $X_{\rm A}^{\rm neq}(t)$, and $U_{\rm A}^{\rm neq}(t)$ and $\Xi_{\rm A}^{\rm neq}(t)$ were evaluated according to the procedure described in Sec. III.3 using the TDL transformations Eqs. (16) and (17). The results are depicted as non–equilibrium $x(t)$$X_{\rm A}^{\rm neq}(t)$ and $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams, which are closed because the cycle is steady and because $T(t)$, $S_{\rm A}^{\rm neq}(t)$, $x(t)$, and $X_{\rm A}^{\rm neq}(t)$ are state variables.
We first present the $x(t)$$X_{\rm A}^{\rm neq}(t)$ diagrams for weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) system–bath coupling strengths in Fig. 3. The area enclosed by each diagram corresponds to positive work when evolving in the counterclockwise direction over time, whereas it corresponds to negative work when evolving clockwise. The work performed in one cycle is presented in Table 4. In each diagram in Fig. 3, the red and blue horizontal lines appear because there is a time delay in the change of $X_{\rm A}^{\rm neq}(t)$ with respect to $x(t)$ for the heat bath to take effect.
As shown in Fig. 3 and Table 4, the larger the SB coupling, the smaller the quantum effect, because it is suppressed by relaxation.Ikeda and Tanimura (2019) Thus, when the coupling is large, dissipation becomes dominant, and the system becomes a refrigerator (or damper). This is because the time scales of fluctuations and dissipation are different: in the non–equilibrium case, dissipation dominates when the SB coupling is large, whereas in the quasi-static case, fluctuations and dissipation are balanced regardless of the coupling strength. This is a distinct difference from the Carnot case; the efficiency reaches a maximum in the intermediate SB coupling region where $A$ is neither large nor small.KT20JCP2
Next, we present the $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams in Fig. 4. The area enclosed by a clockwise curve represents the difference between extensive heat and waste heat, $Q_{\rm A}^{ext}-Q_{\rm A}^{\rm wst}$, per cycle. In the quasi-static case, the area agrees with the extensive heat per cycle because $Q^{\rm wst}$ becomes zero. In the thermostatic processes (green and orange curves), when the SB coupling is small, there is a time delay in the change of $S_{\rm A}^{\rm neq}(t)$ with respect to $T(t)$. This occurs because it takes time for the system to be excited when thermal fluctuations are small, whereas the delay is almost negligible when the SB coupling is strong.
Figure 5: (a) Waste heat current $dQ^{\rm wst}(t)/dt$, (b) enthalpy, and (c) internal energy as functions of $t$. The blue, green, and red curves represent the weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling cases, respectively. The solid and dashed curves represent the quantum and classical results, respectively. The plots of waste heat currents have been time-averaged over 10 steps each using the raw data.
In Fig. 5, we show the (time-averaged) waste heat current, enthalpy, and internal energy as functions of $t$. As we show in Sec. II, the waste heat current is always negative, and it becomes large in the isothermal case for larger SB coupling because of the strong dissipation, whereas it becomes small in the thermostatic processes for the case of weak SB coupling because fluctuations are suppressed.
Using the graph of the waste heat current, we can improve the cycle efficiency. For instance, in the case of strong coupling, because a large waste heat current occurs in the isothermal processes, by increasing the duration of the isothermal processes, we can reduce the waste heat in the cycle.
For weak SB coupling, both the enthalpy and internal energy in the classical case agree with those in the quantum case, whereas for strong SB coupling, they disagree, owing to the quantum coherence between the system and bath (bathentanglement).Tanimura (2020)
IV Conclusions
The virtue of thermodynamics lies in its ability to describe, in a system-independent manner, macroscopic thermal phenomena resulting from complex microscopic interactions as changes in thermodynamic potentials described as interrelated intensive and extended variables through Legendre transformations.
We have developed the laws of thermodynamics as applied to work in a system-independent manner, based on the principle of DL minimum work. Subsequently, we have developed a non–equilibrium thermodynamic theory that describes non–equilibrium thermodynamic potential in terms of time-dependent intensive and extensive variables, along with waste heat.
Thus, the non–equilibrium Massieu–Planck and Helmholtz–Gibbs potentials are obtained in time-derivative form in terms of non–equilibrium extensive and intensive variables, which are state variables along with waste heat, which is not a state variable. Our results are summarized in Tables 1 and 2 and are consistent with traditional thermodynamics in a quasi-static case, as presented in Tables 5 and 6.
These results have been validated for quantum and classical Stirling engines through numerical simulations based on the thermodynamic SB model, which can describe both isothermal and thermostatic processes. An optimization algorithm has been used to evaluate non–equilibrium thermodynamic potentials. Work diagrams in non–equilibrium regimes have been presented and analyzed.
Although the time evolution of the system in the non–equilibrium regime is model-specific, it is possible to test our theory even in real systems through the use of optimization algorithms. By designing processes to reduce waste heat, it may be possible to develop efficient heat engines even for non–equilibrium processes.
In engineering, the effective energy from non-equilibrium to equilibrium is referred to as exergy.Rant (1956); Rosen and Dincer (2001) The non–equilibrium thermodynamic potentials we have introduced can be regarded as a generalization of it. The present theory provides a methodology to systematically evaluate and improve it, which should be useful from the perspective of Sustainable Development Goal (SDGs).
Acknowledgments
The authors thank Yoshi Oono for his critical comments on the definition and terminology of thermodynamic variables, particularly for extensive variables. Y.T. was supported by JSPS KAKENHI (Grant No. B21H01884). S.K. was supported by JST fellowship, the establishment of university fellowships towards the creation of science technology innovation (Grant No. JPMJFS2123), and by Grant-in-Aid for JSPS Fellows (Grant No. 24KJ1373).
Author declarations
Conflict of Interest
The authors have no conflicts to disclose.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Quasi-static thermodynamic potentials
Table 5: Total differential expressions for the quasi-static (qst.) entropic potentials as functions of the intensive variables $\beta^{\rm qst}(t)$ and $\tilde{x}^{\rm qst}(t)$ and the extensive variables $U_{\rm A}^{\rm qst}(t)$ and ${X}_{\rm A}^{\rm qst}(t)$. Entropy has two definitions, depending on whether the work variable is intensive or extensive. Of these DL entropies, the commonly used one, which we call Massieu entropy (M-entropy), involves only extensive variables and is denoted by $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$, whereas the less widely used one, which we call Planck entropy (P-entropy), is denoted by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$. Whereas the enthalpy $H_{\rm A}^{\rm qst}(t)$ was chosen as the natural variable in Ref. Koyanagi and Tanimura, 2024a, here we chose the internal energy $U_{\rm A}^{\rm qst}$ instead. Each potential is related to the others by the Legendre transformations shown in the final column.
Qst. Potential Differential form Natural var. Legendre transformation
Massieu $d\Phi_{\rm A}^{\rm qst}=-U_{\rm A}^{\rm qst}d\beta^{\rm qst}-\tilde{x}^{\rm qst}d{X}_{\rm A}^{\rm qst}$ $\beta^{\rm qst},{X}_{\rm A}^{\rm qst}$cdots
Planck $d\Xi_{\rm A}^{\rm qst}=-U_{\rm A}^{\rm qst}d\beta^{\rm qst}+{X}_{\rm A}^{\rm qst}d\tilde{x}^{\rm qst}$ $\beta^{\rm qst},\tilde{x}^{\rm qst}$ $\Xi_{\rm A}^{\rm qst}=\Phi_{\rm A}^{\rm qst}+\tilde{x}^{\rm qst}{X}_{\rm A}^{\rm qst}$
M-Entropy $d\Lambda_{\rm A}^{\rm qst}=\beta^{\rm qst}dU_{\rm A}^{\rm qst}-\tilde{x}^{\rm qst}dX_{\rm A}^{\rm qst}$ $U_{\rm A}^{\rm qst},X_{\rm A}^{\rm qst}$ $\Lambda_{\rm A}^{\rm qst}=\Xi_{\rm A}^{\rm qst}+\beta^{\rm qst}U^{\rm qst}_{\rm A}$
P-Entropy $d\Gamma_{\rm A}^{\rm qst}=\beta^{\rm qst}dU_{\rm A}^{\rm qst}+X_{\rm A}^{\rm qst}d\tilde{x}^{\rm qst}$ $U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}$ $\Gamma_{\rm A}^{\rm qst}=\Phi_{\rm A}^{\rm qst}+\beta^{\rm qst}U^{\rm qst}_{\rm A}$
We introduce the DL intensive heat defined as
$\frac{d\tilde{Q}_{\rm A}^{int}(t)}{dt}\equiv\beta(t)\frac{dU_{\rm A}(t)}{dt}+X_{\rm A}(t)\frac{d\tilde{x}(t)}{dt}$ (45)
which satisfies the Legendre transformation
$\frac{d\tilde{Q}_{\rm A}^{int}(t)}{dt}=\frac{d\tilde{Q}^{ext}_{\rm A}(t)}{dt}+\frac{d}{dt}\left[\tilde{x}(t)X_{\rm A}(t)\right].$ (46)
The extensive and intensive work and heat $\tilde{W}_{\rm A}^{ext}(t)$, $\tilde{W}_{\rm A}^{int}(t)$, $\tilde{Q}_{\rm A}^{ext}(t)$, and $\tilde{Q}_{\rm A}^{int}(t)$ are interrelated through the Legendre transformations (6), (7), and (46). Therefore, the inequalities (8), (9),
$\tilde{W}_{\rm A}^{int}\geq-\Delta\Xi_{\rm A}^{\rm qst},$ (47)
and
$\tilde{Q}_{\rm A}^{int}\leq\Delta\Gamma_{\rm A}^{\rm qst},$ (48)
where $\Gamma_{\rm A}^{\rm qst}$ is the Planck entropy (P-entropy), are all expressions of the second law of thermodynamics.
The total differential forms of the Massieu–Planck potentials are presented in Table 5.
The Helmholtz–Gibbs potentials can be obtained form the Massieu–Planck potentials using the definitions $F^{\rm qst}_{\rm A}(t)=-\Phi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$ and $G^{\rm qst}_{\rm A}(t)=-\Xi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$. From Eq. (6), we have
$F_{\rm A}^{\rm qst}(t)=G_{\rm A}^{\rm qst}(t)+x^{\rm qst}(t)X_{\rm A}^{\rm qst}(t).$ (49)
Accordingly, from Eq. (7), we have
$U_{\rm A}^{\rm qst}(t)=F_{\rm A}^{\rm qst}(t)+T^{\rm qst}(t)S_{\rm A}^{\rm qst}(t),$ (50)
where we have used $d\beta^{\rm qst}(t)/dt=-(1/k_{\rm B}[T^{\rm qst}(t)]^{2})dT^{\rm qst}(t)/dt$ and $S_{\rm A}^{\rm qst}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm qst}(t)$.
The total differential forms of the Helmholtz–Gibbs potentials are presented in Table 6.
Table 6: Total differential expressions for the quasi-static (qst.) thermodynamic potentials as functions of the intensive variables $T^{\rm qst}(t)$ and $x^{\rm qst}(t)$ and the extensive variables $S_{\rm A}^{\rm qst}(t)$ and ${X}_{\rm A}^{\rm qst}(t)$, which are related through the Legendre transformations shown in the final column.Koyanagi and Tanimura (2024a)
Qst. potential Differential form Natural var. Legendre transformation
Helmholtz $dF_{\rm A}^{\rm qst}=-S_{\rm A}^{\rm qst}dT^{\rm qst}+x^{\rm qst}dX_{\rm A}^{\rm qst}$ $T^{\rm qst},X_{\rm A}^{\rm qst}$cdots
Gibbs $dG_{\rm A}^{\rm qst}=-S_{\rm A}^{\rm qst}dT^{\rm qst}-X_{\rm A}^{\rm qst}dx^{\rm qst}$ $T^{\rm qst},x^{\rm qst}$ $G_{\rm A}^{\rm qst}=F_{\rm A}^{\rm qst}-x^{\rm qst}X_{\rm A}^{\rm qst}$
Internal $dU_{\rm A}^{\rm qst}=T^{\rm qst}dS_{\rm A}^{\rm qst}+x^{\rm qst}dX_{\rm A}^{\rm qst}$ $S_{\rm A}^{\rm qst},X_{\rm A}^{\rm qst}$ $U_{\rm A}^{\rm qst}=F_{\rm A}^{\rm qst}+T^{\rm qst}S_{\rm A}^{\rm qst}$
Enthalpy $dH_{\rm A}^{\rm qst}=T^{\rm qst}dS_{\rm A}^{\rm qst}-X_{\rm A}^{\rm qst}dx^{\rm qst}$ $S_{\rm A}^{\rm qst},x^{\rm qst}$ $H^{\rm qst}_{\rm A}=G_{\rm A}^{\rm qst}+T^{\rm qst}S_{\rm A}^{\rm qst}$
Appendix B DL non–equilibrium minimum work principle
Figure 6: Schematic to derive Eq. (51). The blue circles 1 and 2 represent the equilibrium states, and the red square A represents the non–equilibrium state. The quasi-static work performed for $1\rightarrow 2$ (red dashed line) is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow 2}^{\rm min}$, while the non–equilibrium work from 1 to A (dark blue curve) is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$. When $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$ is fixed, the value of $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ depends on the path (green and light blue curves). From the second law of thermodynamics, we have $(W_{\rm A}^{ext})_{1\rightarrow A}+(\tilde{W}_{\rm A}^{ext})_{\rm A\rightarrow 2}\geq-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}$. This indicates a lower bound of work for ${A\rightarrow 2}$ denoted by $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ (light blue curve) and given by Eq. (51).
Here, we derive fundamental equations to develop the DL non–equilibrium minimum work principle. Consider two equilibrium states 1 and 2 that are connected by a quasi-static process (see Fig. 6). From the principle of DL minimum work [Eq. (8)], the work done in this process is equivalent to the difference in Massieu potentials: $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow 2}^{\rm min}=-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}\equiv(\Phi_{\rm A}^{\rm qst})_{1}-(\Phi_{\rm A}^{\rm qst})_{2}$, where ${n\rightarrow n^{\prime}}$ represents the transition from any state $n$ to $n^{\prime}$. Separately, we consider the non–equilibrium state $A$ and introduce the non–equilibrium process $1\rightarrow A$ (dark blue curve in Fig. 6), and any process from $A$ to 2 (green and light blue curves). For each of these processes, the work done is denoted by $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$ and $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$, respectively. For fixed $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$, the value of $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ changes depending on the path.
From the second law of thermodynamics [Eq. (8)], we then have $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}+(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}>-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}.$ This indicates that $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}$ has a lower bound for fixed $(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$; otherwise the inequality is violated, for example, as $-\infty>-(\Delta\Phi_{\rm A}^{\rm qst})_{1\rightarrow 2}-(\tilde{W}_{\rm A}^{ext})_{1\rightarrow A}$. Using the lower bound of the work, we introduce the non–equilibrium Massieu potential $\Phi_{\rm A}^{\rm neq}$ defined as
$(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}=-[(\Phi_{\rm A}^{\rm qst})_{2}-(\Phi_{\rm A}^{\rm neq})_{A}].$ (51)
Figure 7: Schematic to derive the principle of DL non–equilibrium minimum work expressed by Eq. (53). We consider two non–equilibrium states $A$ and $B$ (red square) that are connected by a non–equilibrium process described by the red arrow for $A\rightarrow B$. The light blue curves $A\rightarrow 2$ and $B\rightarrow 2$ represent non–equilibrium-to-equilibrium minimum work paths whose work is denoted by $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ and $(\tilde{W}_{\rm A}^{ext})_{B\rightarrow 2}^{\rm min}$, respectively.
Introducing a second non–equilibrium state $B$ on the pathway $A\rightarrow 2$ as depicted in Fig. 7, we now discuss the non–equilibrium transition ${A\rightarrow B}$ (red arrow). We consider the non–equilibrium-to-equilibrium minimum work from $B$ to 2 (light blue curve), expressed as
$(\tilde{W}_{\rm A}^{ext})_{B\rightarrow 2}^{\rm min}=-[(\Phi_{\rm A}^{\rm qst})_{2}-(\Phi_{\rm A}^{\rm neq})_{B}].$ (52)
From the inequality $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}\geq-[(\Phi_{\rm A}^{\rm qst})_{2}-(\Phi_{\rm A}^{\rm neq})_{A}]$ and Eq. (52), we have
$(\tilde{W}_{\rm A}^{ext})_{A\rightarrow B}\geq-(\Delta\Phi_{\rm A}^{\rm neq})_{A\rightarrow B},$ (53)
where $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow B}\equiv(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}-(\tilde{W}_{\rm A}^{ext})_{B\rightarrow 2}^{\rm min}$. Thus, equality holds in (53) when $B$ in on the minimal pathway $(A\rightarrow 2)^{\rm min}$.
Because a non–equilibrium process changes as a function of time, it is more convenient to use the time-derivative form of the above. Thus, for the process from $A$ to $B$ at times $t$ and $t+dt$, where $dt$ is an infinitesimal time, we have the following inequality:
$\frac{d\tilde{W}_{\rm A}^{ext}(t)}{dt}\geq-\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}.$ (54)
This inequality is an extension of the principle of minimum work to a non–equilibrium regime.
Figure 8: Schematic showing that the non–equilibrium Massieu potential can be evaluated from any equilibrium state 3 instead of 2. To illustrate this, we consider the non–equilibrium-to-equilibrium path $A\rightarrow 3$, where we have introduced a third equilibrium state denoted by 3 (blue circle). The work between the two equilibrium states is denoted by $(W_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}$ and $(W_{\rm A}^{ext})_{3\rightarrow 2}^{\rm qst}$, respectively.
Although we have derived Eqs. (53) and (54) for the specific equilibrium state 2, actually the value $(\Phi_{\rm A}^{\rm neq})_{A}$ is the same for any quasi-equilibrium state along $\beta^{\rm qst}(t)$ and $x^{\rm qst}(t)$. To illustrate this, we introduce a third equilibrium state 3 and consider the transition $A\rightarrow 3$ (see Fig. 8). Because $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}+(\tilde{W}_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}\geq(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}$ for $A\rightarrow 2\rightarrow 3$ and $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}+(\tilde{W}_{\rm A}^{ext})_{3\rightarrow 2}^{\rm qst}\geq(\tilde{W}_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}$ for $A\rightarrow 2\rightarrow 3$, we have
$(W_{\rm A}^{ext})_{A\rightarrow 2}^{\rm min}+(\Phi_{\rm A}^{\rm neq})_{2}=(W_{\rm A}^{ext})_{A\rightarrow 3}^{\rm min}+(\Phi_{\rm A}^{\rm neq})_{3},$ (55)
where we have used $(W_{\rm A}^{ext})_{2\rightarrow 3}^{\rm qst}=(\Phi_{\rm A}^{\rm qst})_{3}-(\Phi_{\rm A}^{\rm qst})_{2}$. From Eq. (51), this indicates that the non–equilibrium Massieu potential is independent of the choice of the equilibrium state 2.
Because the other potentials are interrelated through TDL transformations, we have
$\frac{d\tilde{W}_{\rm A}^{int}(t)}{dt}\geq-\frac{d\Xi_{\rm A}^{\rm neq}(t)}{dt},$ (56)
$\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}\leq\frac{d\Lambda_{\rm A}^{\rm neq}(t)}{dt},$ (57)
and
$\frac{d\tilde{Q}_{\rm A}^{int}(t)}{dt}\leq\frac{d\Gamma_{\rm A}^{\rm neq}(t)}{dt}.$ (58)
For an isothermal case, from the definitions $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta$ and $G^{\rm neq}_{\rm A}(t)=-\Xi_{\rm A}^{\rm neq}(t)/\beta$, we have
$\frac{dW_{\rm A}^{ext}(t)}{dt}\geq\frac{dF_{\rm A}^{\rm neq}(t)}{dt}$ (59)
and
$\frac{dW_{\rm A}^{int}(t)}{dt}\geq\frac{dG_{\rm A}^{\rm neq}(t)}{dt}.$ (60)
However, unlike in the Massieu–Planck case, the equality in Eqs. (59) and (60) cannot be obtained in general because the minimum DL work path in Fig. 6 is evaulaed by optimizing both $\beta(t)$ and $x(t)$, while $\beta$ has been fixed to introduce the Helmholtz energy as $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta$.
The above discussion indicates that by including the thermostatic process, we may develop an efficient heat engine under non–equilibrium conditions.
Appendix C Convexity of non–equilibrium thermodynamic potentials as functions of extensive variables
Table 7: Inequalities holding between the non–equilibrium state at time $t_{1}$ and the equilibrium state at time $t_{2}$, together with the conditions under which these inequalities hold. From these inequalities, we find that, for example, for a given inverse temperature and electric field, the non–equilibrium Planck potential takes its maximum value when the state is in equilibrium.
neq Potentials Convexity and Concavity Condition
Massieu $(1-\lambda)(\Phi_{\rm A}^{\rm neq})_{A}+\lambda(\Phi_{\rm A}^{\rm neq})_{B}\leq(\Phi_{\rm A}^{\rm neq})_{\lambda}$ $\beta_{A}=\beta_{B}=\beta_{\lambda}$
Planck $(1-\lambda)(\Xi_{\rm A}^{\rm neq})_{A}+\lambda(\Xi_{\rm A}^{\rm neq})_{B}\leq(\Xi_{\rm A}^{\rm neq})_{\lambda}$ $\beta_{A}=\beta_{B}=\beta_{\lambda}$ and $\tilde{x}_{A}=\tilde{x}_{B}=\tilde{x}_{\lambda}$
Helmholtz $(1-\lambda)(F_{\rm A}^{\rm neq})_{A}+\lambda(F_{\rm A}^{\rm neq})_{B}\geq(F_{\rm A}^{\rm neq})_{\lambda}$ $\beta_{A}=\beta_{B}=\beta_{\lambda}$
Gibbs $(1-\lambda)(G_{\rm A}^{\rm neq})_{A}+\lambda(G_{\rm A}^{\rm neq})_{B}\geq(G_{\rm A}^{\rm neq})_{\lambda}$ $\beta_{A}=\beta_{B}=\beta_{\lambda}$ and $x_{A}=x_{B}=x_{\lambda}$
Figure 9: Schematic for deriving convexity relations. The red squares represent the non–equilibrium states. The arrows are the relaxation paths to equilibrium state 2 described by the time-evolution operator $\hat{U}_{\rm A}(t,t_{0})$: The blue arrow is the DL minimum work path from state $\lambda$.
To obtain stable thermodynamic properties, the thermodynamic potential must be a convex function of the work variables. Even in the non–equilibrium case, we find that convexity holds under certain conditions. Here, we show this in the case of the SB model. Because other potentials can be discussed similarly, we limit our discussion here to the case of a non–equilibrium Massieu potential.
We consider three non–equilibrium states $A$, $B$, and $\lambda$ ($\lambda\in[0,1]$) with the same initial inverse temperature $\beta(t_{0})$ at time $t_{0}$. We then assume that the state $\lambda$ satisfies the relation
$(W_{\rm A})_{\lambda}=(1-\lambda)(W_{\rm A})_{A}+\lambda(W_{\rm A})_{B},$ (61)
[…]
$({X}_{\mathrm{A}})_{\lambda}(t)=(1-\lambda)({X}_{\mathrm{A}})_{A}(t)+\lambda({X}_{\mathrm{A}})_{B}(t)$ (62)
and
$(U_{\mathrm{A}})_{\lambda}(t)=(1-\lambda)(U_{\mathrm{A}})_{A}(t)+\lambda(U_{\mathrm{A}})_{B}(t),$ (63)
where we have defined $(X_{\mathrm{A}})_{\alpha}(t)=\mathrm{tr}_{\rm tot}\{X_{\rm A}(q)\hat{U}_{\rm tot}(t,t_{0})(W_{\rm tot})_{\alpha}\}$ and $(U_{\mathrm{A}})_{\alpha}(t)=\mathrm{tr}_{\rm tot}\{[p^{2}/2m+U(q)]\hat{U}_{\rm tot}(t,t_{0})(W_{\rm tot})_{\alpha}\}$ for $\alpha=A,B,$ and $\lambda$. Thus, from Eq. (2), the DL extensive work $(W_{\rm A}^{ext})_{\alpha}(t,t_{0})$ ($\alpha=A,B,$ and $\lambda$), performed between times $t_{0}$ and $t$ also satisfies the equality
$(W_{\rm A}^{ext})_{\lambda}(t,t_{0})=(1-\lambda)(W_{\rm A}^{ext})_{A}(t,t_{0})+\lambda(W_{\rm A}^{ext})_{B}(t,t_{0}).$ (64)
The DL minimum work path from the state $\lambda$ to the equilibrium state 2 is described by $\hat{U}_{\rm A}(t,t_{0})$ (see Fig. 9). This indicates that when $t$ is large, the intensive variables must be $\beta(t)\rightarrow\beta_{2}$ and ${x}(t)\rightarrow x_{2}$, where $\beta_{2}$ and $x_{2}$ are the values at the equilibrium state 2. Therefore, the processes starting from $A$ and $B$ also relax to the equilibrium state 2 as $t\rightarrow\infty$. Then, from Eq. (64), we obtain
$(W_{\rm A}^{ext})_{\lambda\rightarrow 2}^{\rm min}=(1-\lambda)(W_{\rm A}^{ext})_{A\rightarrow 2}+\lambda(W_{\rm A}^{ext})_{B\rightarrow 2}.$ (65)
Using the inequality $(W_{\rm A}^{ext})_{\alpha\rightarrow 2}\geq(W_{\rm A}^{ext})_{\alpha\rightarrow 2}^{\rm min}$ for the states $\alpha=A$ and $B$, and Eq. (51), we obtain the inequality to prove convexity as
$(1-\lambda)(\Phi_{\rm A}^{\rm neq})_{A}+\lambda(\Phi_{\rm A}^{\rm neq})_{B}\leq(\Phi_{\rm A}^{\rm neq})_{\lambda},$ (66)
where $(\Phi_{\rm A}^{\rm neq})_{\alpha}$ is the non–equilibrium Massieu potential in the state $\alpha$ ($\alpha=A$, $B$, and $\lambda$). When $A$ and $B$ are in equilibrium states, we obtain the convexity relation in the quasi-static case for fixed inverse temperature using the inequality for the Massieu potential in Table 8 as Koyanagi and Tanimura (2024a)
$(1-\lambda)(\Phi_{\rm A}^{\rm qst})_{A}+\lambda(\Phi_{\rm A}^{\rm qst})_{B}\leq(\Phi_{\rm A}^{\rm qst})_{\lambda}.$ (67)
We summarize the convexity and concavity properties for the other non–equilibrium thermodynamic potential in Table 7.
Appendix D Inequality between non–equilibrium and quasi-static thermodynamic potentials
Table 8: Inequalities for Massieu–Planck potentials holding between the non–equilibrium state $A$ at time $t_{1}$ and the equilibrium state 2 at time $t_{2}$, together with the conditions under which these inequalities hold. From these inequalities, we find that, for example, for a given inverse temperature and intensive variable, the non–equilibrium Planck potential takes its maximum value when $A$ is in the equilibrium state.
neq Potentials Inequalities Condition
Massieu $\Phi_{\rm A}^{\rm qst}(t_{2})\geq\Phi_{\rm A}^{\rm neq}(t_{1})$ $\beta(t_{1})=\beta(t_{2})$ and $X_{\rm A}(t_{1})=X_{\rm A}(t_{2})$
Planck $\quad\Xi^{\rm qst}_{\rm A}(t_{2})\geq\Xi_{\rm A}^{\rm neq}(t_{1})\quad$ $\beta(t_{1})=\beta(t_{2})$ and $\tilde{x}(t_{1})=\tilde{x}(t_{2})$
C–Entropy $\Lambda_{\rm A}^{\rm qst}(t_{2})\geq\Lambda_{\rm A}^{\rm neq}(t_{1})$ $U_{\rm A}(t_{1})=U_{\rm A}(t_{2})$ and $X_{\rm A}(t_{1})=X_{\rm A}(t_{2})$
B–Entropy $\Gamma_{\rm A}^{\rm qst}(t_{2})\geq\Gamma_{\rm A}^{\rm neq}(t_{1})$ $U_{\rm A}(t_{1})=U_{\rm A}(t_{2})$ and $\tilde{x}(t_{1})=\tilde{x}(t_{2})$
Table 9: Inequalities for Helmholtz–Gibbs potentials holding between the non–equilibrium state at time $t_{1}$ and the equilibrium state at time $t_{2}$, together with the conditions under which these inequalities hold.
neq Potentials Inequalities Condition
Helmholtz $F_{\rm A}^{\rm qst}(t_{2})\leq F_{\rm A}^{\rm neq}(t_{1})$ $T(t_{1})=T(t_{2})$ and $X_{\rm A}(t_{1})=X_{\rm A}(t_{2})$
Gibbs $G_{\rm A}^{\rm qst}(t_{2})\leq G_{\rm A}^{\rm neq}(t_{1})$ $T(t_{1})=T(t_{2})$ and $x(t_{1})=x(t_{2})$
Internal Energy $U_{\rm A}(t_{2})\leq U_{\rm A}(t_{1})$ $S_{\rm A}^{\rm neq}(t_{1})=S_{\rm A}^{\rm qst}(t_{2})$ and $X_{\rm A}(t_{1})=X_{\rm A}(t_{2})$
Enthalpy $H_{\rm A}(t_{2})\leq H_{\rm A}(t_{1})$ $S_{\rm A}^{\rm neq}(t_{1})=S_{\rm A}^{\rm qst}(t_{2})$ and $x(t_{1})=x(t_{2})$
Under given conditions, the non–equilibrium DL thermodynamic potentials are smaller than the quasi-static ones. Here, we derive the inequalities between the quasi-static and non–equilibrium thermodynamic potentials.
We consider a non–equilibrium-to-equilibrium transition $A\rightarrow 2$ from time $t_{1}$ to $t_{2}$ described by the inverse temperature $\beta(t)$ and the extensive variable $X_{\rm A}^{\rm neq}(t)$. The DL extensive work in this process is evaluated as
[…]
$=\tilde{x}_{2}(X_{\rm A}^{\rm neq}(t_{2})-X_{\rm A}^{\rm neq}(t_{1})),$ (68)
where we have assumed that $\beta(t_{1})=\beta_{2}$ and $X_{\rm A}^{\rm neq}(t_{1})=(X_{\rm A}^{\rm qst})_{2}$, and that $\beta_{2}=\beta(t)$ and $\tilde{x}_{2}=\tilde{x}(t)$ for time $t>t_{1}$, to make the subsystem relax to the equilibrium state 2 at time $t_{2}$. Because $X_{\rm A}^{\rm neq}(t_{2})=(X_{\rm A}^{\rm qst})_{2}=X_{\rm A}^{\rm neq}(t_{1})$, the right-hand side of Eq. (D) vanishes. Thus, the net DL extensive work is zero, and so the DL minimum work principle reduces to $0\geq(\Phi_{\rm A}^{\rm neq})_{A\rightarrow 2}$, from which it follows that $\Phi_{\rm A}^{\rm qst}(t_{2})\geq\Phi_{\rm A}^{\rm neq}(t_{1})$.
The inequalities for the other DL thermodynamic potentials are obtained in the same manner. We summarize the results in Tables 8 and 9
Figure 10: Difference between the quasi-static and non–equilibrium Planck potential in the quantum and classical cases (solid and dashed curves, respectively) for one cycle of the Stirling engine. The blue, green, and red curves represent the cases of weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling, respectively.
In Fig. 10, to illustrate the inequalities presented in Table 8, we plot the difference between the quasi-static and non–equilibrium Planck potential, $\Delta\Xi_{\rm A}(t)=\Xi_{\rm A}^{\rm qst}[\beta(t),\tilde{x}(t)]-\Xi_{\rm A}^{\rm neq}(t)$, for one cycle of the Stirling engine described in Sec. III.4. As can be seen, $\Delta\Xi_{\rm A}(t)$ is always positive and satisfies the inequalities in Table 8.